OpenStudy (anonymous):

2√2(√2 - √10) + √3(√3 + √15)

OpenStudy (jiteshmeghwal9):

\[4-4\sqrt{5}+3+3\sqrt{5}\]\[7-\sqrt{5}\]

OpenStudy (jiteshmeghwal9):

right @mayankdevnani ??

OpenStudy (mayankdevnani):

absolutely...

OpenStudy (anonymous):

That's what I got, I just wanted to check my work, thanks.